### v1

@@ 1 Introduction
In this work, we propose ProtoMIL, an approach to automatically discover pathology concepts and use them to train an inherently interpretable MIL framework. To achieve this, we leverage recent advances in mechanistic interpretability, employing a sparse autoencoder (SAE) [3, 14, 11] to learn pathology concepts automatically. These discovered concepts can be inspected through representative patches from a probing set, allowing users to understand the concepts better and identify diagnostically irrelevant ones. ProtoMIL is then trained using the discovered concepts, with the option to eliminate unwanted concepts through human intervention. While both concept discovery and human-model interaction have been explored in computer vision [10, 17, 16, 14] and natural language processing [3, 5], they are underexplored in computational pathology, especially for WSI analysis. To our knowledge, this is the first work in computational pathology to demonstrate the effectiveness of automatic concept discovery and interventions on concepts within an MIL framework.
@@ 4 Discussion and Conclusion
This paper proposes a novel, inherently interpretable MIL model for WSI analysis. Our experimental results show that SAE effectively learns sparse, human-interpretable pathology concepts. Using the activations of the learned concepts, we train an inherently interpretable ProtoMIL model, which achieves comparable classification performance with state-of-the-art MIL models while providing easy-to-understand explanations. A key advantage of our approach is that it allows model interventions to eliminate dependence on spurious signals, potentially improving the model’s reliability. In future work, we plan to explore strategies to improve SAE and mitigate the presence of polysemantic neurons.
5 Acknowledgment
This work was funded by the Deutsche Forschungsgemeinschaft (DFG) – EXC number 2064/1 – Project number 390727645 and the Carl Zeiss Foundation in the project “Certification and Foundations of Safe Machine Learning Systems in Healthcare”. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Susu Sun.

### v2

@@ 1 Introduction
In this work, we propose ProtoMIL, an approach to automatically discover pathology concepts and use them to train an inherently interpretable MIL framework. To achieve this, we leverage recent advances in mechanistic interpretability, employing a sparse autoencoder (SAE) [3, 15, 12] to learn pathology concepts automatically. These discovered concepts can be inspected through representative patches from a probing set, allowing users to understand the concepts better and identify diagnostically irrelevant ones. ProtoMIL is then trained using the discovered concepts, with the option to eliminate unwanted concepts through human intervention. While both concept discovery and human-model interaction have been explored in computer vision [11, 18, 17, 15] and natural language processing [3, 6], they are underexplored in computational pathology, especially for WSI analysis. A related approach is ProtoMixer [5], which clusters image embeddings to derive prototypes and applies an MLP-Mixer architecture for cancer subtyping, primarily aiming to reduce WSI size and improve computational efficiency. However, its reliance on cluster centroids may limit interpretability, as the prototypes do not necessarily correspond to semantically meaningful pathology concepts. To our knowledge, our framework is the first in computational pathology to demonstrate the effectiveness of automatic concept discovery and interventions on concepts within an MIL framework.
@@ 4 Discussion and Conclusion
This paper proposes a novel, inherently interpretable MIL model for WSI analysis. Our experimental results show that SAE effectively learns sparse, human-interpretable pathology concepts. Using the activations of the learned concepts, we train an inherently interpretable ProtoMIL model, which achieves comparable classification performance with state-of-the-art MIL models while providing easy-to-understand explanations. A key advantage of our approach is that it allows model interventions to eliminate dependence on spurious signals, potentially improving the model’s reliability. As we have observed that a certain number of neurons in the SAE remain polysemantic, in future work, we plan to conduct a more thorough investigation of the learned concept space and develop methods to further improve the SAE and reduce polysemantic representations. Additionally, we plan to explore the feasibility of developing a more generalized SAE capable of learning transferable concepts across diverse tissue categories and cancer types, thereby providing interpretable concept representations for a wide range of downstream tasks.
{credits}
4.0.1 Acknowledgements
Funded by the Deutsche Forschungsgemeinschaft (DFG) – EXC number 2064/1 – Project number 390727645, the Carl Zeiss Foundation in the project “Certification and Foundations of Safe Machine Learning Systems in Healthcare”, the Dutch Research Council, the European Research Council, the European Union Horizon Europe Program, and the Hanarth Fund. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Susu Sun.